OpenStudy (anonymous):

A rectangle has length 4x –1 and a width 2x + 1 -a) Show how to find the perimeter of the rectangle b) Simplify the expression. -a) Write an algebraic expression for the area of the rectangle b) Simplify the expression - If the area of another rectangle is x2 + 5x + 4, find algebraic expressions representing possible lengths and widths of the rectangle.

OpenStudy (anonymous):

perimeter=2(l+b)=2( (4x-1)+(2x+1) )

OpenStudy (anonymous):

perimeter=2(4x+2x+0)=2(6x)=12x

OpenStudy (anonymous):

area = 8x^2+2x-1

OpenStudy (anonymous):

is that for the last one?

OpenStudy (anonymous):

when we factorize x^2+5x+4 we get (x+4)(x+1) so the possible lengths are (x+4) and (x+1)

OpenStudy (anonymous):

no thats the b part of ques 2..i accedently deleted its a part

OpenStudy (anonymous):

its a part is area=(4x-1)(2x+1)

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

my pleasure:D